Again, this is from my nephews Home School course, and it stumps me. Thank you in advance for your help. "One number is 11 more than twice another number. If the sum of the numbers is twice their difference, find the numbers."
Answer by scott8148(6628)

let x="first number", so 2x+11="second number"

the sum of the numbers is twice their difference ___ x+2x+11=2(2x+11-x)

3x+11=4x+22-2x ___ 3x+11=2x+22